OpenStudy (anonymous):

Max rented a motorbike at $465 for 5 days. If he rents the same motorbike for a week, he has to pay a total rent of $625. Write an equation in the standard form to represent the total rent (y) that Max has to pay for renting the motorbike for x days. (4 points) Part B: Write the equation obtained in Part A using function notation. (2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (anonymous):

part A: Write an equation in the standard form to represent the total rent (y) that Max has to pay for renting the motorbike for x days. (4 points)

OpenStudy (anonymous):

@peachpi I already have part A: though.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@phi

OpenStudy (phi):

what did you get for A?

OpenStudy (anonymous):

part A: y=65+80(x) plug in 5. 65+80(5) y=65+400. y=465. when he rents the bike for five days, he pays 465. so if he rents it for a week he would be renting it for7 days (x)days. so y=65+80(7) 65+560. y=625.

OpenStudy (phi):

that is the correct equation, but I do not know what the rest is. If you want to show your work, start by saying you have these two pairs: (5,465) and (7,625) and find the equation of a line through two points.

OpenStudy (anonymous):

the rest would be what he has to pay in the week and I had to show how I got to 625

OpenStudy (phi):

first find the slope = change in y divided by change in x = (625-465)/(7-5) = 160/2 = 80 so y= 80x + b using (5,465) in that to find "b" 465= 80*5 + b 465= 400 + b b= 465-400 = 65 so the equation is y= 80x+65

OpenStudy (phi):

part B, use "function notation" I think they want you to write f(x) = 80x+65

OpenStudy (anonymous):

ok. and so for part C?

OpenStudy (phi):

part C. I would say , plot two points on the line, and "connect the dots" x axis is days y axis is rental in dollars

OpenStudy (phi):

to make it fit on the graph better, you might want each "step" in the y-direction be say 20 or 40 dollars (rather than 1)

OpenStudy (anonymous):

ok thank you.